RE: Shift report 20030116 08:00-16:00

From: Ian Bearden (bearden@nbi.dk)
Date: Fri Jan 17 2003 - 10:39:02 EST

  • Next message: Ian Bearden: "RE: TOFW, RICH"
    On 16. Jan JJG wrote in the shift report:
    
    6. First 16 channels of H1 slat distribution alternate in
    abnormally in yield. TRIGGER BOARD problem??? EXPERT MUST INVESTIGATE.
    
    7. H2: TDCTOP ch 31 is dead.
    On 17. Jan, I write:
    I am starting the ADC calibrations for the TOF detectors, and have seen this
    problem.
    In H1, slat 1 is OK, slat 2 has very few counts.  This pattern repeats up
    slat 15 OK, slat 16 very count, and then things look OK.  Has anyone looked
    into this?
    It dosn't seem to be a problem related to pulse height, as the ADC peak is
    roughly where it should be (though with far too few counts!).
    This is a critical problem (assuming we get beam and HPSS at the same time)
    so
    please take a look.
    
    Thanks,
    Ian
     --------------------------------------------------------------------
    | I.G. Bearden                                                       |
    | Niels Bohr Institute Tb 3   email:      bearden@nbi.dk             |
    | Blegdamsvej 17              phone:      (+45) 35 32 53 23          |
    | København Ø                   FAX:      (+45) 31 42 10 16          |
    | Danmark                       Www:      http://www.nbi.dk/~bearden |
     --------------------------------------------------------------------
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri Jan 17 2003 - 10:39:16 EST